OpenStudy (anonymous):

\[\sin ^{3}\alpha=3\sin \alpha-4\sin ^{3}\alpha\]

OpenStudy (anonymous):

identities

OpenStudy (asnaseer):

you just need to find the value for \(\alpha\) - correct?

OpenStudy (anonymous):

no, i need to prove they are equal

OpenStudy (asnaseer):

to simplify, just set \(x=\sin\alpha\) to get:\[x^3=3x-4x^3\]solve for x and then you can find \(\alpha\)

OpenStudy (mertsj):

It's not an equation asnaseer. It's and identity

OpenStudy (asnaseer):

how can it be an identity if it is only true for a certain set of values for \(\alpha\)?

OpenStudy (mertsj):

It's true for all values of alpha.

OpenStudy (anonymous):

\[\sin (\alpha )=\frac{e^{i \alpha }-e^{-i \alpha }}{2 i}\] \[\left(\frac{e^{i \alpha }-e^{-i \alpha }}{2 i}\right)^3\]= \[\frac{3}{4} \frac{\left(i e^{-i \alpha }- i e^{i \alpha }\right)}{2}-\frac{1}{4}\frac{\left(i e^{-3 i \alpha }- i e^{3 i \alpha }\right)}{2}\] \[\frac{3}{4}(\text{Sin} \alpha )-\frac{1}{4}\text{Sin} 3\alpha\]

OpenStudy (mertsj):

That is the problem. To show that the two sides are the same thing.

OpenStudy (asnaseer):

take \(\alpha=90\), we get: 1 = 3 - 4 which is not true.

OpenStudy (asnaseer):

@tomaasc - maybe you have stated the identity incorrectly? - please double check.

OpenStudy (anonymous):

asnaseer is right, it's not an identity.

OpenStudy (anonymous):

It is \[\sin[x]^3=\frac{3}{4}(\text{Sin} \alpha )-\frac{1}{4}\text{Sin} 3\alpha\] not what you wrote

OpenStudy (mertsj):

Then he copied the problem wrong because he is trying to prove identities.

OpenStudy (asnaseer):

@imranmeah91 has stated a correct identity - is that the one you wanted to prove @tomaasc?

OpenStudy (anonymous):

maybe the left side is 3sin

OpenStudy (mertsj):

Well, it's your problem. Don't you know what it is/

OpenStudy (anonymous):

yes 3 Sin[x]^3 is what you wrote